OpenStudy (anonymous):

Select the equation of the line that passes through the point (2, –1) and has slope 3 in point-slope form. A. (y + 1) = 3(x – 2) B. (x – 2) = 3(y + 1) C. (y – 1) = 3(x + 2) D. (x + 2) = 3(y – 1)

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

Equation line in slope popint form is given as; y-y1=m(x-x1) here x1=2 , y1=-1 and slope m=3 Now plugging these values in the above eqution we find y-(-1)=3(x-2) i.e. y+1=3(x-2) Thus option A. (y + 1) = 3(x – 2) is the correct answer. @TakeA.I.M.